OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (anonymous):

first find the slope m=(-3-6)/(0-3)=3 now use slope intercept form y=mx+C y=3x+C use x=3 and y=6 to find C.

OpenStudy (anonymous):

Isn't that MX+b..

OpenStudy (anonymous):

its up to u use C or b its just a constant.

OpenStudy (anonymous):

Oh okay.. That makes sense I have had MB+X pounded in my head when I first took math is all.

OpenStudy (anonymous):

that,s right dear